OpenStudy (anonymous):

How much MnO2(s) should be added to excess HCl(aq) to obtain 305 mL of Cl2(g) at 25 °C and 705 Torr?

OpenStudy (cuanchi):

1. write down the equation and balance it 2.Convert the mL of Cl2 to moles ( n= RT/PV) 3. calculate the moles of MnO2 according to the stoiquiometry of the reaction in 1 4 calculate the mass of MnO2 ( m = n x MM )

OpenStudy (anonymous):

That's what I did I got 1.1649g

OpenStudy (anonymous):

it's calling it wrong

OpenStudy (cuanchi):

sorry, n= PV/RT I typed wrong before

OpenStudy (anonymous):

well wouldn't that be the same if i was solving for n?

OpenStudy (cuanchi):

how many moles of Cl2 did you get?

OpenStudy (anonymous):

.0134

OpenStudy (cuanchi):

grate!, what is the stoiquiometry of the reaction and the molecular mass of the MnO2?

OpenStudy (anonymous):

well the number of moles of MnO2 is the same, right? so i took that and multiplied it by the MW of MnO2

OpenStudy (cuanchi):

yes, what is the answer?

OpenStudy (anonymous):

1.1382 which isn't right

OpenStudy (cuanchi):

I got 1.04g

OpenStudy (anonymous):

how?!

OpenStudy (cuanchi):

n= (0.928 x 0.305) /(0.082 x 298) = 0.0116 moles 0.0116 moles x 86.94 g/mol = 1.007 g

OpenStudy (anonymous):

Thank you!

OpenStudy (cuanchi):

you welcome!!